Qs=18000+0.2p

Qd=2400-0.1p


1
Expert's answer
2021-03-09T15:17:21-0500

Market equilibrium occurs when "Q_d=Q_s":


"24000-0.1P_E=18000+0.2P_E,""6000=0.3P_E,""P_E=\$20000."

Finally, we can find the equilibrium quantity:


"Q_E=24000-0.1\cdot\$20000=22000."